AnimeGhoul8863:

Juliana has created the function f(x) = 3x+2 over 4 to represent the cost of texting on her current plan, where x represents the number of texts. Juliana discovers that, using the inverse function to solve for x = 24, she can predict how many texts she can use for $24. Explain to Juliana how to accomplish this, using complete sentences.

Elsa213:

Magic is the way. e.e

AnimeGhoul8863:

really?

Elsa213:

Ofc Ofc

AnimeGhoul8863:

@Shadow

Shadow:

Do you know how to get the inverse of a function?

AnimeGhoul8863:

sorta im not strong at it no but i can try

Shadow:

Well you know it or you don't. Do you know the first step to getting the inverse of a function?

AnimeGhoul8863:

First, replace f(x) with y

AnimeGhoul8863:

then switch the x and y's

Shadow:

Yes. f(x) and y are interchangeable. f(x) stands for the function of x. So y is the function of x. You don't need to switch them but it makes your life easier. Then you switch the positions of x and y. Then you solve for the variable that you switched to the right side. So it would go like: \[f(x) = \frac{ 3x + 2 }{ 4 }\] \[y = \frac{ 3x + 2 }{ 4 } \] \[x = \frac{ 3y + 2 }{ 4 } \] \[4x = 3y + 2\] \[4x - 2 = 3y \] \[\frac{ 4x - 2 }{ 3} = y \] Cosmetic change, cause this is how we usually visualize equations. \[y = \frac{ 4x - 2 }{ 3 }\]

AnimeGhoul8863:

ok showing Juliana the equation is explaining to her how to accomplish it?

Shadow:

Well they want it in words, 'complete sentences.' I expressed that in my above response.

AnimeGhoul8863:

so i would write down Juliana in order to accomplish your goal the first thing you need to do is take f(x) and change it to y because f(x) and y are interchangeable. Next we talk x and y and switch them so 3x would turn into 3y and instead of y= it would be x=. Doing it this way makes life easier. Then you solve for the variable that you switched to the right side.

AnimeGhoul8863:

^is that correct

Shadow:

I think your lesson has a typo. x represents the # of texts, yet they note $24 as being a value for x. I am sure they mean y.

Shadow:

Try to use your own words, not mine. Teachers can usually tell when you copy and paste as they have plenty of samples of your work.

Shadow:

So you would put 24 in for a value of y, as we found the inverse function and isolated y. Then you can solve for the number of texts that you can send with $24.

AnimeGhoul8863:

the only thing of yours i used was "f(x) and y are interchangeable." and "Then you solve for the variable that you switched to the right side." thats it the other 90% is my own words

AnimeGhoul8863:

wait you just confused me alittle bit what do you mean

AnimeGhoul8863:

were do we add the 24?

Shadow:

Okay, let's go back to the beginning. What are we trying to do?

AnimeGhoul8863:

first we switch f(x) with y then switch y with X

Shadow:

Okay that's the first step of obtaining the inverse function. But why do they want us to do that?

AnimeGhoul8863:

they want us to inverse because it makes the equation easier to solve?

Shadow:

Anime I am giving you my time to help you understand. I could easily type out the answer and you'd get an A but you would never learn and would likely fail any tests. If you don't want to be present for my help, that's okay. I have an essay to write. But I'd appreciate if you stayed out of chat when I'm waiting for a response from you.

Shadow:

Is that fair?

AnimeGhoul8863:

what are you talking about shadow im giving my time to its not like i wanna be here im waiting on your responses so i talk in the chat i see you doing the same thing during helping people it not my fault i didnt reply that exact moment you press post

Shadow:

That's because I already have stuff posted for them to go through and respond to.

AnimeGhoul8863:

Im doing my best to learn this as fast as i can and give the answers im suppose to give just give me a break this is difficult for me im not genius like you people

Shadow:

I already understand this material.

Shadow:

Lets move on. Read the question again. Why are they having you do all this math? What do they want you to solve for?

Shadow:

The latter question is a hint.

AnimeGhoul8863:

they want me to solve for the amount of text that can be sent with only $24

Shadow:

Correct. So what do x and y represent?

AnimeGhoul8863:

the amount of text being sent

AnimeGhoul8863:

and the money being used for it

Shadow:

x is the first or second one

AnimeGhoul8863:

first x= amount of text y= Amount of money

Shadow:

Correct. So does that answer your question of where you should add the 24?

AnimeGhoul8863:

Juliana in order to accomplish your goal the first thing you need to do is take f(x) and change it to y because f(x) and y are interchangeable. Next we take x and y and switch them so 3x would turn into 3y and instead of y= it would be x=. Doing it this way makes life easier. You also need to remember x= the amount of text being sent and y= the amount of money you can spend. Then you solve for the variable that you switched to the right side.

AnimeGhoul8863:

^good?

Shadow:

Sure. Just know that the practice of verbatim (repeating someone else's words) in note taking or homework doesn't instill concepts as strongly. When you take their words and put it into your own, you remember it better. If you'd like I could show you the study, but the disparity in test scores is massive when it comes to conceptual questions. But sure, that works.

AnimeGhoul8863:

i used 2 things you said not the whole thing 90% is my words

Shadow:

So after you solve for the variable on the right side, what do you get/have to do?

AnimeGhoul8863:

simplify?

Shadow:

So we are at \[y = \frac{ 4x - 2 }{ 3 } \] Input 24 in for y, solve for x. You get the # of texts you can send with $24

AnimeGhoul8863:

ummmmmmmmmmm so 24=4x-2 over 3

Shadow:

Yes

AnimeGhoul8863:

switch the sides \[\frac{ 4x-2 }{ 3 }=24\]

AnimeGhoul8863:

Then...............Multiply by 3 on both sides? \[\frac{ 3(4x-2) }{ 3 }=24 \times3\]

Shadow:

Mhm

Shadow:

PEMDAS

mikewwe13:

Shadow don't mean to interrupt could you help me after this question ?

AnimeGhoul8863:

thennnnnn..............4x-2 which = 2?

AnimeGhoul8863:

right

Shadow:

Add two to both sides, then divide by 3.

Shadow:

Don't know what 4x-2 which = 2 means

AnimeGhoul8863:

PEMDAS () first so (4x-2) so i put 4x-2 = 2

Shadow:

How does 4x - 2 = 2

AnimeGhoul8863:

I dont know i was just thinking 4-2=2 which would make it 3*2 over 3

mikewwe13:

x = 1

AnimeGhoul8863:

does it = 1/2

Shadow:

Okay so you are correct that you do parentheses first when it comes to PEMDAS but you can't do anything with (4x -2). \[y = \frac{ 4x - 2 }{ 3 }\] You want to substitute 24 in for y so we can get the # of texts we can send with $24. \[24 = \frac{ 4x - 2 }{ 3 }\] Multiply both sides by 3 \[72 = 4x - 2\] Add two to both sides \[74 = 4x \] Divide both sides by 4 \[x = 18.5\] What do you think the correct answer is?

AnimeGhoul8863:

18.5 text per 24 dollars she spends

Shadow:

You can't send half a text

AnimeGhoul8863:

18.5 dollars for the text she sends monthly

Shadow:

What variable did we just solve for

AnimeGhoul8863:

x

Shadow:

Which is the # of texts you can send in a month

AnimeGhoul8863:

texts messages I SAID THAT BUT U SAID U CANT SEND A HALF A MESSAGE

Shadow:

Yeah but you can send 18

AnimeGhoul8863:

ok i got it i think thx shadow i have a few more but help mike out

AnimeGhoul8863:

Hes been patiently waiting

Shadow:

mhm